# Sharpening Kubernetes Audit Logs with Context Awareness

### v1

@@ 8 Conclusion
We presented an innovative approach for the analysis of K8s audit logs, called K8NTEXT. K8NTEXT employs a combination of heuristics and ML techniques to automatically aggregate and contextualize audit logs, making them more comprehensible and reducing both their complexity and the volume of their analysis. K8NTEXT is efficient, extensible, and scales well with the size of the cluster, enabling real-time inspection of K8s audit logs. The output of K8NTEXT is space-efficient and is both human- and machine-readable, allowing for further analysis and integration with other tools while preserving the context of the actions performed in the cluster.

### v2

@@ 8 Conclusion
We presented an innovative approach for the analysis of K8s audit logs, called K8NTEXT. K8NTEXT parses, aggregates, and contextualizes raw audit logs using a ML model and a clustering algorithm, yielding a compact representation of the activities within a K8s cluster. K8NTEXT is efficient and scalable, making it suitable for real-time inspection of K8s audit logs, both by humans and machines.
As future work, we plan to explore techniques for automatic generation and labeling of K8s scenarios, as well as methods for incremental learning to adapt the model to new scenarios with minimal human intervention. Additionally, we aim to investigate the integration of K8NTEXT with other security tools to enhance monitoring and detection capabilities within K8s environments.